A company's costs, in millions of dollars, are given by the equation, C=x^2-3x-27, where x is the number of items sold, in thousands. How many items must be sold for the costs to be zero?

Answer by lwsshak3(11628)

A company's costs, in millions of dollars, are given by the equation, C=x^2-3x-27, where x is the number of items sold, in thousands. How many items must be sold for the costs to be zero?
***
C=x^2-3x-27
x^2-3x-27=0
solve for x by quadratic formula:
x≈6.908
How many items must be sold for the costs to be zero? approximately 6908